OpenStudy (anonymous):

“The captain of a rafting boat takes his crew down river for 6 hours, covering a total of 56 miles. When he realizes he left one of his passengers behind, he is forced to paddle back upriver for 8 miles. The trip back up took him 3 hours.” A student used the following system of equations to model the situation given above. Which of the statements is true about the system of equations the student used? b = the speed of the boat r = the speed of the currentA. The system of equations the student used is correct. B. The top equation in the system is incorrect. C. The bottom equation in the

OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. x intercept = (9, 0); y intercept = (0, -6) x intercept = (6, 0); y intercept = (0, -9) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (anonymous):

...system are correct

OpenStudy (anonymous):

A. The system of equations the student used is correct. B. The top equation in the system is incorrect. C. The bottom equation in the system is incorrect. D. Both equations in the system are incorrect.

OpenStudy (the_fizicx99):

Make a system? Using elimination or sub?

OpenStudy (anonymous):

elimination

OpenStudy (the_fizicx99):

The 8's are kinda confusing lol xD 8b -8r = 3 6b + 6r = 56 I normally use this method:  Then you would eliminate r and just add the b's and the 18 + 448 it would be:  Would you say 4.8541666666666666666666666666667 rounded be 4.9 or 5.0